WAC Notes July 26 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes July 25 2019 << >> WAC Notes July 27 2019

July 26 2019

Daily Meeting BlueJeans

Production Run List

Start Slug 14, IHWP In, FFB on always

    • Since we are using bcm_dg_ds, the width in the bcm asymmetry is 2X of what we had with bcm_an_ds
    • We used the same, bcm_dg_ds, bcm in slug13 also, where the Aq width was ~1.5X of what we had with bcm_an_ds
    • SAM3 started saturating since run3640
    • Should we think about changing saturation limit for this SAM?
    • If we compare with 70uA good run from previous slug, yield of SAMs on Right Hand Side has increased while on the other side decreased
  • Good - run3649ADDED from suspicious list
    • Short run with many beam trips
    • Caryn says "attempting to go to production but having some FFB issues"
  • Good - run3650ADDED from suspicious list
    • BPM11 saturated for a long time
  • Good - run3651ADDED from suspicious list
    • BPM11 has saturated couple of times
  • Good - run3652
  • Good - run3653
  • Good - run3654
  • Good - run3655
  • Good - run3656
  • Good - run3657
  • Good - run3658
  • Good - run3659
  • Good - run3664
    • Regressed detector widths are ~5ppm higher than previous runs->might be due to the jump in current from 70uA to 60uA
  • Good - run3666
    • mini-run 8 had too large widths in regressed detector asymmetry
    • Quad problem, applied cut->Looks good

Non-Production Runs and Bad Runs

  • Test run3637
  • Test run3638
  • Test run3639
  • Test run3642
  • Test run3643
    • Musson injects 10uA fake signal into digital system after new chassis added
    • All the runs upto here were under slug13 conditions
  • Test run3644
    • Target Trouble Shooting
  • Test run3647
    • Beam current ~20uA. No event passed event cuts.
  • Bad run3648
    • Attempting to production but some Feedback issue
    • Quite big Aq mean, very short run ~11k events
    • Only 4k events have passed event cuts
  • No Beam - run3665
  • Test - run3667
  • Test - run3668

Pedestal/Calibrations Updates

Issues

Other Status Updates